A rectangular pot of ground measuring 12 meters by 20 meters is surrounded by a sidewalk of uniform width. The area of the sidewalk is 68 square meters. Find the width of the sidewalk.
Answer by mananth(16946)

Let the width of sidewalk be x
..
Length of ground =20
Width = 12

Area = 20*12= 240M^2
Area of sidewalk =68M^2
Length of ground + sidewalk =20+2x
width of ground + sidewalk = 12+2x
...
(20+2x)(12+2x) -240=68

240+40x +24x+4X^2-240=68

4X^2+64x-68=0
/4
x^2+16x-17=0
x^2+17x-x+17=0
x(x+17)-1(x+17)=0
(x+17)(x-1)=0
x= 1 the width of side walk.
...
m.ananth@hotmail.ca